find Ssey R R is a triangular region in x-y plane with vertices (-2, 2), (0,0), (2, 2)
Answer #1

Let I= sse ey R where is R Ky plane with triangular region in vertices (+2, 2), 100), (2,2). Y 2,2 (-2,29 0,0) Х -y si Shadedy=- and eqin of y-Y Is Thy - Yi (x-x1) In Region R, limit > Y-0=(2-0) (-0) (2-0) - egn of line Joining Prod and (2, 2) Pling)> I- 0 egy evady Syeddy Put 4= 4 =) du dy gy dy I du > I= ody L & Se Use special integral de la duz -2 1,=- FC8,-43) Resubsti